V=10 sin 2π100t


1
Expert's answer
2021-03-22T08:36:55-0400

Let the velocity of the body change according to the law "v=10\sin(2\pi100t)" . How its acceleration changes?


"a=\frac{dv}{dt}=10\cos(2\pi100t)\cdot2\pi100=2000\pi\cos(2\pi100t)"


"a_{max}=2000\pi"